OpenStudy (anonymous):

6x + 2y = 0 4x + y = -1 solve using substitution

OpenStudy (amoodarya):

6x + 2y = 0 4x + y = -1 ---> y=-1-4x put y in first equation 6x+2(-1-4x)=0 6x-2-8x=0 -2x=+2 x=-1 now put x in one of equation

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

x + 3y = -5 2x - y = 18

OpenStudy (anonymous):

work out what x is equal to in the first equation, the top one.

OpenStudy (anonymous):

it's \[x = -5-3y\]

OpenStudy (anonymous):

x+3y=-5 =3y=-5+x= =y=-5/3+x/3

OpenStudy (anonymous):

2x-y=18 _y=18-2x soooo replace in the first equation y by the answer above x+3(18-2x)=-5 x+54-6x=-5 x-6x=-5-54 -5x=-59 x=-59/-5 x=59/5

OpenStudy (anonymous):

\[2(-5-3y)-y=18\] \[-10-6y-y=18\] -7y =28 y= -4

OpenStudy (anonymous):

@julian9922 x= 7 and y = -4